# PO-CKAN:Physics Informed Deep Operator Kolmogorov Arnold Networks with Chunk Rational Structure

### v1

@@ Abstract
We propose PO-CKAN, a physics-informed deep operator framework based on Chunkwise Rational Kolmogorov–Arnold Networks (KANs), for approximating the solution operators of partial differential equations. This framework leverages a Deep Operator Network (DeepONet) architecture that incorporates Chunkwise Rational Kolmogorov–Arnold Network (CKAN) sub-networks for enhanced function approximation. The principles of Physics-Informed Neural Networks (PINNs) are integrated into the operator learning framework to enforce physical consistency. This design enables the efficient learning of physically consistent spatio-temporal solution operators and allows for rapid prediction for parametric time-dependent PDEs with varying inputs (e.g., parameters, initial/boundary conditions) after training. Validated on challenging benchmark problems, PO-CKAN demonstrates accurate operator learning with results closely matching high-fidelity solutions. PO-CKAN adopts a DeepONet-style branch–trunk architecture with its sub-networks instantiated as rational KAN modules, and enforces physical consistency via a PDE residual (PINN-style) loss. On Burgers’ equation with $\nu=0.01$, PO-CKAN reduces the mean relative $L^{2}$ error by approximately 48% compared to PI-DeepONet, and achieves competitive accuracy on the Eikonal and diffusion–reaction benchmarks.
@@ 1 Introduction
The Physics-Informed DeepONet (PI-DeepONet) [29] is an emerging machine learning framework designed to efficiently solve entire families of parameterized partial differential equations (PDEs) [35, 36, 20, 37, 42, 26]. Unlike traditional Physics-Informed Neural Networks (PINNs) [43], which solve a single problem, PI-DeepONet is designed to approximate the underlying solution operator. This operator constitutes a universal mapping from any given input function (e.g., boundary conditions, forcing terms) to its corresponding solution function [40, 13, 7, 43, 48, 58, 24, 47]. It achieves this through a unique ”Branch–Trunk” architecture and ensures that its learning process adheres to fundamental physical laws by incorporating the governing equation residuals into its loss function. Its most prominent advantage is its ”train once, predict many” efficiency, which drastically reduces computational expense in scenarios requiring large-scale repeated solutions. A promising frontier in this area involves enhancing the PI-DeepONet framework by replacing its MLP components with a more expressive architecture, an approach that has already shown success [12, 29, 34].
Kolmogorov-Arnold Networks (KANs) [32, 25, 46, 31] present a compelling alternative to conventional Multilayer Perceptrons (MLPs) [23] and are thus well-suited for this application. KANs, which are theoretically grounded in the Kolmogorov-Arnold Representation Theorem [39], introduce a significant architectural departure from MLPs. Unlike MLPs, which pair fixed activation functions on the nodes with learnable weights and biases [5, 50], KANs place learnable activation functions, originally parameterized as B-splines [52, 19], on the edges of the network. This fundamental difference in architecture yields notable benefits, such as improved model interpretability and more robust performance on tasks characterized by noisy data or requiring continual learning [56, 11, 18, 30, 53, 22, 41].
@@
Importantly, this theoretical power has been validated in the practical application of solving differential equations. For instance, KANs have been successfully integrated into PINNs to solve the Poisson equation [57], and a KAN-based operator network, DeepOKAN [2], was introduced to simulate sinusoidal wave solvers and address orthotropic elasticity problems using Radial Basis Functions (RBFs). Both studies confirmed that KANs significantly outperform traditional MLP architectures. However, despite this demonstrated superiority, a critical flaw obstructs their widespread adoption. The total number of learnable parameters in KANs grows quadratically with network width. This severe scalability bottleneck renders a naive integration into the DeepONet framework computationally impractical for most operator learning tasks [14, 8, 15], creating a significant gap between the architecture’s theoretical potential and its practical application.
To overcome this fundamental challenge, we introduce our primary technical innovation: the Chunkwise Rational KAN (CKAN). This architecture is specifically designed for parameter efficiency. It employs a chunk-wise parameter sharing mechanism, where chunks of connections share a common activation function, while individual edges retain unique scalar weights. To further enhance performance, we leverage the power of rational functions for these activations. Foundational work has established that rational functions, which have the form $R(x)=P(x)/Q(x)$, are highly effective approximators, particularly for functions with complex features such as singularities or steep gradients [6, 10, 49, 27]. Integrating such learnable rational functions into the KAN structure presents a compelling direction for creating more powerful and efficient neural networks [3, 38]. Our CKAN design actualizes this concept. This novel combination of chunking and rational activations breaks the quadratic scaling bottleneck, making the expressive power of KANs computationally feasible for large-scale operator learning.
By integrating our efficient CKAN architecture into the physics-informed DeepONet structure, we construct our final model: the Physics-Informed Deep Operator Chunk-rational KAN (PO-CKAN). This framework inherits the operator learning paradigm from its DeepONet foundation while leveraging the enhanced and scalable expressiveness of our CKAN architecture to achieve superior accuracy and greater computational efficiency. The primary contributions of this study are:
@@
    We introduce an efficient rational activation function, a lightweight alternative to splines that enhances computational efficiency while maintaining high expressive power.
@@
    We propose the Physics-Informed Operator CKAN (PO-CKAN) by integrating our scalable CKAN into the DeepONet architecture, which to our knowledge marks the first successful incorporation of KAN-like expressivity into physics-informed operator learning in a scalable manner.
@@
    We demonstrate the effectiveness and versatility of the PO-CKAN framework across a comprehensive suite of four benchmarks. Our model successfully tackles diverse problem classes—from time-evolution systems (e.g., Burgers’ equation) to boundary-value problems (e.g., Eikonal equation)—and consistently achieves a superior accuracy-efficiency trade-off compared to traditional DeepONet baselines.
@@ 2.1.1 Deep Operator Network (DeepONet)
The DeepONet framework is designed to approximate solution operators for PDEs. It employs a distinctive “Branch–Trunk” architecture to learn the mapping from an input function, $u(y)$, to a solution function, $s(x)$. The core components are:
@@
    Branch Network: Encodes the input function $u(y)$ into a $p$-dimensional feature representation, $b=[b_{1},\dots,b_{p}]\in\mathbb{R}^{p}$.
@@
    Trunk Network: Processes the spatio-temporal coordinates, $x$, to learn a set of $p$ coordinate-dependent basis functions, $t(x)=[t_{1}(x),\dots,t_{p}(x)]$.
The final solution, $s(x)$, is then reconstructed as the dot product between the outputs of the two networks:
@@ 2.1.2 Kolmogorov–Arnold Network (KAN)
A foundational result in function approximation theory, the Kolmogorov–Arnold representation theorem, offers a powerful alternative to the conventional paradigms of neural network design [21]. The theorem reveals that the apparent complexity of multivariate continuous functions is reducible; that is, any such function on a bounded domain can be expressed as a finite composition of single-variable functions linked by addition. For a specific smooth function $f:[0,1]^{n}\to\mathbb{R}$, the theorem guarantees that the following representation holds:
@@
In this formulation, both the inner functions $\phi_{q,p}:[0,1]\to\mathbb{R}$ and the outer functions $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are univariate. This remarkable theoretical statement suggests that simple addition is the only multivariate operation required. However, while theoretically profound, the functions guaranteed by the original theorem can be non-smooth and pathological, which has historically limited its direct application and hindered attempts to develop practical, learnable models based on it.
The Kolmogorov–Arnold Network (KAN) [32] provides a learnable framework for the theorem by modeling its univariate functions with B-splines [17, 45, 44, 1, 9]. Fundamentally differing from MLPs, KANs employ learnable activation functions on the edges of the network rather than fixed activations on the nodes.
A KAN is constructed as a sequence of layers defined by its shape $n_{0},n_{1},\dots,n_{L}]$. A matrix of learnable activation functions connects each layer to the next. The output of a neuron in a given layer is the sum of the outputs from all of its incoming functions. This process is described by the equation:
@@
The full network is therefore a composition of these function matrices, $\mathbf{\Phi}_{l}$:
@@
This structure is fundamentally different from a Multilayer Perceptron (MLP). An MLP alternates between linear transformations, $\mathbf{W}$, and a single, fixed non-linear function, $\sigma$:
@@
The key distinction, therefore, is that KANs embed both linear transformations and nonlinearities into the learnable function matrices $\mathbf{\Phi}$.
This design gives KANs superior expressive power and interpretability. However, it also creates a major challenge. The activation functions are typically parameterized using B-splines. This B-spline parameterization yields a large number of parameters and incurs a substantial computational cost. Consequently, the standard KAN design does not scale well. A more efficient and parsimonious architecture is therefore needed for large-scale applications.
@@ 2.2.1 Overall Architecture
The PO-CKAN architecture inherits the “Branch-Trunk” structure from DeepONet, as shown in Figure 1. The model has two main components: a branch net and a trunk net. The branch net processes the input function $u(y)$, while the trunk net processes the coordinates $x$. Both nets are built with our novel CKAN layers, which replace traditional MLPs. This substitution is designed to capture complex functional relationships while maintaining computational tractability.
@@ 2.2.2 CKAN Layer
As previously noted, the KAN architecture’s high expressiveness is coupled with major scalability challenges: the high computational cost of B-spline activations and the prohibitive parameter count. To address these challenges, we introduce the Chunk-rational KAN (CKAN), a novel layer design featuring two key innovations.
@@
To enhance the computational efficiency and expressive power of the KAN layer, we move beyond the traditional B-spline basis and instead employ rational functions. A rational function, which is simply the ratio of two polynomials, offers superior approximation properties for complex behaviors and is highly efficient to compute [27, 10]. Specifically, we parameterize each edge function $\phi(x)$ as a weighted rational function of the form $w\frac{P(x)}{Q(x)}$, where the numerator $P(x)$ and denominator $Q(x)$ are polynomials of degrees $m$ and $n$, respectively.
$\phi(x)=wF(x)=w\cdot\frac{P(x)}{Q(x)}=w\cdot\frac{a_{0}+a_{1}x+\cdots+a_{m}x^{m}}{b_{0}+b_{1}x+\cdots+b_{n}x^{n}}$ (3)
However, the standard form in Eq. 3 is susceptible to numerical instability from poles, which occur where the denominator $Q(x)$ approaches zero. To ensure stable training while retaining high expressiveness, we adopt the Enhanced Rational Unit (ERU) as our basis function, $F(x)$ [51]. The ERU takes the form:
$F(x)=ax+b+\sum^{\frac{n}{2}}_{k=1}\frac{c_{k}x+d_{k}}{(x-e_{k})^{2}+f_{k}^{2}+\epsilon}$ (4)
This formulation, with its always-positive denominator, effectively prevents the occurrence of poles. This choice is further justified by its significant computational advantages. As shown in Table 1, the ERU is substantially more efficient to compute than the original B-spline activations.
Name FLOPs
B-Spline (G=3, K=3) 204
Rational (m=5, n=4) 46
ERU ($n=4$) 19
Table 1: Comparison of FLOPs for different functions. Using the Enhanced Rational Unit reduces FLOPs by approximately 10.7x compared to the B-Spline function.
Chunk KAN
To tackle the challenge of excessive parameters, we introduce a Chunk-wise Parameter Sharing mechanism to create the CKAN. The core concept is to share the base parameters of a rational function across a chunk of connections rather than learning a unique function for each input-output pair. Figure 2 shows the difference between the CKAN, KAN, and a standard MLP. Specifically, for a CKAN layer mapping a $d_{\text{in}}$-dimensional input to a $d_{\text{out}}$-dimensional output:
@@
Fig. 2: Comparison of our CKAN (with a 2x2 chunk configuration) with a standard KAN and MLP. While a KAN has a unique function on each input-output pair, a CKAN shares a single base function across a chunk of edges.
Suppose $i$ indexes the input channel. With $c\times c$ chunks, the output of the $j$-th neuron in a CKAN layer is given by:
$\text{CKAN}(x)_{j}=\sum_{i=1}^{d_{\text{in}}}w_{ij}\cdot F_{\lfloor i/d_{\text{in},c}\rfloor,\lfloor j/d_{\text{out},c}\rfloor}(x_{i})$ (5)
where $d_{\text{in},c}=d_{\text{in}}/c$ and $d_{\text{out},c}=d_{\text{out}}/c$ are the chunk sizes for the input and output dimensions, respectively, and $(\lfloor i/d_{\text{in},c}\rfloor,\lfloor j/d_{\text{out},c}\rfloor)$ is the 2D index of the chunk to which the edge $(i,j)$ belongs.
In the original KAN formulation, $d_{\text{in}}\times d_{\text{out}}$ distinct activation functions are required. By applying our chunking strategy, this number is reduced to only $c\times c$. In addition to parameter reduction, this approach also decreases computational cost: each input chunk evaluates the base function $F$ once, and the result is reused across the same chunk. Conversely, a vanilla KAN computes a separate $\phi_{i,j}$ for every input–output pair, resulting in substantially higher computational complexity.
Example: Consider a layer with $d_{\text{in}}=d_{\text{out}}=50$, B-spline base function degree $K=5$, and grid size $G=50$. In a vanilla KAN, there are $50\times 50=2500$ unique activation functions. Each base function requires $\text{Func FLOPs}\approx 8K=40$ FLOPs, and the total layer FLOPs is approximately $6.7\times 10^{6}$. For a CKAN with $d_{\text{in}}=d_{\text{out}}=50$, rational degree $4$, and a $2\times 2$ chunk configuration, the total FLOPs is roughly $4.4\times 10^{3}$a reduction by three orders of magnitude.
A comparison of parameter counts and computational requirements is summarized in Table 2.
Name No. Params FLOPs
MLP $d_{in}\times d_{out}+d_{out}$ $\text{Func FLOPs}\times d_{out}+2\times(d_{in}\times d_{out})$
KAN $d_{in}\times d_{out}\times(G+K+3)+d_{out}$ $\text{Func FLOPs}\times d_{in}+(d_{in}\times d_{out})\times[9K(G+1.5K)+2G-2.5K+3]$
CKAN $d_{in}\times d_{out}+d_{out}+(2n+2)\times c^{2}$ $(4.5n+1)\times d_{in}\times c+2\times(d_{in}\times d_{out})$
Table 2: Comparison of parameter counts across different models. Here, Func FLOPs refers to the computational cost of the non-linear activation functions. In KAN, $K$ denotes the order and $G$ the number of grid points. For our CKAN, $n$ indicates the rational order, and $c\times c$ specifies the number of chunks. CKAN maintains a parameter count with only a constant overhead compared to a standard MLP, while the KAN parameters grow with $(G+K+3)$.

### v2

@@ Abstract
We present PO-CKAN, a physics-informed deep operator framework based on Chunkwise Rational Kolmogorov–Arnold Networks (CKANs), for approximating the solution operators of partial differential equations. This framework integrates CKAN sub-networks within the Deep Operator Network (DeepONet) architecture, enabling physically consistent and computationally efficient operator learning. A physics-informed residual loss ensures that the learned operator adheres to governing PDE constraints. By embedding PDE-residual losses, PO-CKAN generalizes efficiently across varying parameters, boundary conditions, and initial states. Validated on challenging benchmark problems, PO-CKAN consistently achieves high-fidelity solutions and outperforms PI-DeepONet baselines. PO-CKAN adopts a DeepONet-style branch-trunk architecture with CKAN modules, and enforces physical consistency via a PDE residual (PINN-style) loss. On Burgers’ equation with $\nu=0.01$, PO-CKAN reduces the mean relative $L^{2}$ error by approximately 48% compared to PI-DeepONet, and achieves competitive accuracy on the Eikonal and diffusion-reaction benchmarks.These results highlight PO-CKAN as a scalable and interpretable framework for physics-consistent operator learning
@@ 1 Introduction
The Physics-Informed DeepONet (PI-DeepONet) [27] is an emerging machine learning framework designed to efficiently solve entire families of parameterized partial differential equations (PDEs) [33, 34, 19, 35, 40, 25]. Unlike standard physics-informed neural networks (PINNs), which typically solve individual PDE instances,[41], PI-DeepONet approximates a universal functional mapping from an input function to its corresponding solution function. [38, 13, 7, 41, 46, 56, 23, 45]. It achieves this through a distinctive “branch–trunk” architecture and ensures physical consistency by incorporating governing-equation residuals directly into the loss function. Its key advantage is the “train-once, predict-many” paradigm, which drastically reduces computational cost in settings requiring repeated PDE evaluations.
A continuing research direction seeks to enhance the expressive power and generalization capability of PI-DeepONet by replacing its multilayer-perceptron (MLP) submodules with more advanced operator-learning architectures—an approach that has already yielded promising results. [12, 27, 32].
Kolmogorov-Arnold Networks (KANs) [30, 24, 44, 29] have recently emerged as compelling alternatives to conventional MLPs. [22] and are thus well-suited for this application. Grounded in the Kolmogorov–Arnold Representation Theorem, KANs differ fundamentally by placing learnable activation functions on the network edges rather than using fixed nonlinearities on the nodes.[37, 5, 48] This architectural distinction leads to enhanced interpretability and robustness, particularly in problems with noisy data or continual-learning requirement [54, 11, 18, 28, 50, 21, 39].
@@
Empirical studies have demonstrated that KANs can substantially outperform traditional MLPs in solving differential equations. For instance, KANs have been integrated into PINNs to solve the Poisson equation [55], and the KAN-based operator model DeepOKAN[2], has been applied to wave-propagation and elasticity problems, achieving significant performance gains. However, despite their superior expressiveness, standard KANs suffer from a severe scalability bottleneck: the number of learnable parameters grows quadratically with network width, rendering naive integration into DeepONet computationally impractical for large-scale operator learning tasks [14, 8, 15]. This gap between theoretical potential and practical feasibility motivates the present work.
To address this challenge, we propose the Chunk-wise Rational Kolmogorov–Arnold Network (CKAN)—a parameter-efficient variant that employs chunk-wise parameter sharing and rational activation functions to drastically reduce computational overhead. Rational functions $R(x)=P(x)/Q(x)$ are well-known universal approximators capable of representing functions with singularities or sharp gradients [6, 10, 47, 26]. By embedding such activations into a chunk-shared KAN structure, we obtain a scalable yet expressive operator-learning module.[3, 36].
Integrating this efficient CKAN architecture into the physics-informed DeepONet framework yields the Physics-Informed Deep Operator Chunk-rational KAN (PO-CKAN)—a unified model that couples physical constraints with data-driven operator learning. The major contributions of this study are as follows:
@@
    We introduce an efficient rational activation function, providing a lightweight and numerically stable alternative to B-splines that preserves high expressive power while reducing computational cost.
@@
    We develop the Physics-Informed Operator CKAN (PO-CKAN), integrating the CKAN within the DeepONet architecture to achieve the first scalable fusion of KAN-based expressivity and physics-informed operator learning.
@@
    We demonstrate PO-CKAN’s accuracy and efficiency across four benchmark problems—including Burgers’, Eikonal, fractional, and diffusion–reaction PDEs—showing consistent improvements over PI-DeepONet baselines.
@@ 2.1.1 Deep Operator Network (DeepONet)
The Deep Operator Network (DeepONet) is a neural-operator framework designed to learn mappings between entire functional spaces rather than pointwise variable pairs.Unlike traditional surrogate models that approximate individual mappings between input and output variables, DeepONet directly learns the operator $\mathcal{G}:u(y)\rightarrow s(x)$ that transforms an input function $u(y)$ to its corresponding solution function $s(x)$. It employs a two-stream “branch–trunk” architecture consisting of specialized subnetworks:
@@
    Branch Network: Encodes the input function $u(y)$ into a $p$-dimensional latent representation, $b=[b_{1},\dots,b_{p}]\in\mathbb{R}^{p}$.
@@
    Trunk Network: Processes the spatio-temporal coordinates, $x$, to produce coordinate-dependent basis functions, $t(x)=[t_{1}(x),\dots,t_{p}(x)]$.
The final prediction is expressed as the inner product between branch and trunk features:
@@
This decomposition decouples functional input encoding from spatial representation, yielding a flexible and memory-efficient structure for high-dimensional operator learning.
Once trained, a single DeepONet generalizes to new boundary conditions, forcing terms, or parameter settings without retraining, realizing the hallmark “train-once, predict-many” paradigm of operator learning.
Despite its strong performance, the expressive capacity of DeepONet is bounded by the representational limits of its MLP-based submodules, motivating the development of enhanced architectures—such as the proposed CKAN modules introduced in the next subsection—to better capture complex nonlinear functional dependencies.
@@ 2.1.2 Kolmogorov–Arnold Network (KAN)
The Kolmogorov–Arnold Network (KAN) is a neural architecture grounded in the Kolmogorov–Arnold Representation Theorem[20], which asserts that any multivariate continuous function can be expressed as a finite superposition of univariate functions and addition operations. This theorem provides a rigorous theoretical foundation for constructing neural networks capable of universal function approximation through one-dimensional transformations only.
For a specific smooth function $f:[0,1]^{n}\to\mathbb{R}$, the theorem guarantees that the following representation holds:
@@
where both the inner functions $\phi_{q,p}:[0,1]\to\mathbb{R}$ and the outer functions $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are univariate. This formulation implies that addition is the only multivariate operation required, a remarkably elegant result. Yet, the original theorem’s constructive form involves highly irregular functions, so practical implementation demands smooth, learnable approximations.
To achieve this, the modern KAN parameterizes the univariate functions using smooth basis functions such as B-splines [17, 43, 42, 1, 9]. Unlike conventional MLPs—which alternate between linear transformations and fixed nonlinear activations—KANs place learnable activation functions on the network edges instead of on the nodes.
A KAN with layer structure $n_{0},n_{1},\dots,n_{L}]$. defines a sequence of function-matrix transformations:
@@
so that the overall network $\mathbf{\Phi}_{l}$ is:
@@
By contrast, an MLP uses linear weight matrices $\mathbf{W}$, and a single, fixed non-linear function, $\sigma$:
@@
Thus, KANs integrate both the linear transformation and nonlinearity into learnable edge functions $\mathbf{\Phi}$, endowing them with higher expressivity and interpretability.
However, this expressive power comes at a cost. B-spline parameterization introduces a large number of coefficients and significant computational overhead, leading to quadratic parameter scaling with network width and hampering large-scale deployment.
Therefore, a more compact and computationally efficient variant is essential for practical operator-learning tasks—an issue directly addressed by the proposed Chunk-rational KAN (CKAN) architecture described next.
@@ 2.2.1 Overall Architecture
The Physics-Informed Deep Operator Chunk-rational Kolmogorov–Arnold Network (PO-CKAN) inherits the canonical “branch–trunk” structure of DeepONet, as illustrated in 1. he architecture comprises two complementary components: a branch network that processes the input function $u(y)$, and a trunk network that encodes the spatial–temporal coordinates $x$. Both subnetworks are constructed entirely from our newly developed Chunk-rational KAN (CKAN) layers, which replace the conventional MLP blocks used in standard DeepONet.
This substitution enables the model to capture complex nonlinear functional dependencies while remaining computationally tractable. Each CKAN layer provides a learnable, parameter-efficient transformation that maintains the representational richness of Kolmogorov–Arnold Networks but mitigates their quadratic parameter growth through chunk-wise parameter sharing and rational activation functions.
The overall PO-CKAN framework thus unifies physics-based constraints, operator learning, and scalable architecture design. During training, the model minimizes a composite physics-informed loss function incorporating data fidelity, initial/boundary conditions, and PDE residual terms (as detailed in Section 2.2.3).
Collectively, these innovations enable PO-CKAN to deliver high-fidelity, physics-consistent operator approximations with markedly improved efficiency compared with PI-DeepONet baselines.
@@ 2.2.2 CKAN Layer
The Chunk-rational Kolmogorov–Arnold Network (CKAN) constitutes the core architectural innovation of PO-CKAN, addressing the long-standing scalability limitations of standard KANs.
Traditional KANs exhibit quadratic parameter growth with network width due to the use of edge-specific B-spline activations, leading to high memory cost and computational overhead.
The CKAN layer overcomes these bottlenecks through two complementary mechanisms:
  • 1.
@@
  • 2.
    Chunk-wise Parameter Sharing
Efficient and Stable Rational Activations
To enhance expressiveness while maintaining numerical stability, CKAN replaces the conventional B-spline basis with smooth rational activation functions. [26, 10].
Rational forms are particularly adept at approximating functions that contain sharp gradients, discontinuities, or singularities.
Each edge function is parameterized as:
$\phi(x)=wF(x)=w\cdot\frac{P(x)}{Q(x)}=w\cdot\frac{a_{0}+a_{1}x+\cdots+a_{m}x^{m}}{b_{0}+b_{1}x+\cdots+b_{n}x^{n}},$ (3)
where the numerator $P(x)$ and denominator $Q(x)$ are polynomials of degrees $m$ and $n$, respectively.
To prevent numerical instability from zero denominators, CKAN adopts the Enhanced Rational Unit (ERU) [49], , which guarantees a positive denominator and smooth behavior:
$F(x)=ax+b+\sum^{\frac{n}{2}}_{k=1}\frac{c_{k}x+d_{k}}{(x-e_{k})^{2}+f_{k}^{2}+\epsilon},$ (4)
where $\epsilon>0$ ensures stability.
This design yields significantly lower FLOP counts than the traditional B-spline activation (about $10\times$ reduction) while retaining high expressive power, as summarized in Table 1
Hence, ERUs enable CKAN layers to achieve efficient and numerically robust nonlinearity modeling within a physics-informed operator-learning context.
Model Setting FLOPs
B-Spline G=3,K=3G=3,\ K=3 204
Rational m=5,n=4m=5,\ n=4 46
ERU $n=4$ 19
Table 1: FLOP analysis of various function types. The Enhanced Rational Unit achieves roughly a $10.7\times$ reduction in FLOPs compared with the B-Spline formulation.
Chunkwise Parameter Sharing
The second major improvement, chunk-wise parameter sharing, mitigates the $O(d_{\text{in}}d_{\text{out}})$ growth of activation functions in vanilla KANs.
The core concept is to share the base parameters of a rational function across a chunk of connections rather than learning a unique function for each input-output pair:
@@
Fig. 2: A comparison between our CKAN (configured with a 2×2 chunk) and standard KAN and MLP models. Unlike the conventional KAN, which assigns a distinct function to every input–output connection, CKAN employs one shared base function for each chunk of edges.
Formally, for a CKAN layer mapping $x\in\mathbb{R}^{d}_{\text{in}}$ to $\mathbb{R}^{d}_{\text{out}}$
$\text{CKAN}(x)_{j}=\sum_{i=1}^{d_{\text{in}}}w_{ij}\cdot F_{\lfloor i/d_{\text{in},c}\rfloor,\lfloor j/d_{\text{out},c}\rfloor}(x_{i}),$ (5)
where $d_{\text{in},c}=d_{\text{in}}/c$ and $d_{\text{out},c}=d_{\text{out}}/c$ are the chunk sizes for the input and output dimensions, respectively, and $(\lfloor i/d_{\text{in},c}\rfloor,\lfloor j/d_{\text{out},c}\rfloor)$ identifies the chunk index.
In a standard KAN, $d_{\text{in}}\times d_{\text{out}}$ distinct activations must be evaluated; By CKAN reduces this to only $c\times c$. Consequently, both parameter count and FLOP complexity decrease by several orders of magnitude while maintaining nearly identical inference latency in practice.
Consequently, both parameter count and FLOP complexity decrease by several orders of magnitude while maintaining nearly identical inference latency in practice.
Illustratively, Figure 2 compares KAN, CKAN, and MLP layers: while a standard KAN uses a unique activation per connection, CKAN reuses shared ERU bases within each chunk, dramatically reducing computational overhead.
Table 2 further quantifies this gain, showing that CKAN retains MLP-like parameter counts while achieving KAN-level functional flexibility.
Architecture Parameter Count Estimated FLOPs
MLP $d_{\text{in}}\,d_{\text{out}}+d_{\text{out}}$ $d_{\text{out}}\cdot\text{Func FLOPs}+2\,d_{\text{in}}d_{\text{out}}$
KAN $d_{\text{in}}\,d_{\text{out}}\,(G+K+3)+d_{\text{out}}$ $d_{\text{in}}\cdot\text{Func FLOPs}+d_{\text{in}}d_{\text{out}}\,[\,9K(G+1.5K)+2G-2.5K+3\,]$
CKAN $d_{\text{in}}\,d_{\text{out}}+d_{\text{out}}+(2n+2)c^{2}$ $d_{\text{in}}\cdot\,(4.5n+1)\,c+2\,d_{\text{in}}d_{\text{out}}$
Table 2: Comparison of parameter complexity and computational cost across different neural architectures. Here, Func FLOPs denotes the floating-point operations required to evaluate the non-linear activation function. In KAN, $K$ specifies the base function degree and $G$ the number of grid points. CKAN employs rational functions of order $n$ with a $c\times c$ chunk configuration, resulting in a modest fixed parameter overhead relative to a standard MLP, whereas KAN scales linearly with $(G+K+3)$.